# Towards shutdownable agents via stochastic choice

### v1

@@ 5 Reward functions and agents
DREST reward function. We train agents to be USEFUL and NEUTRAL using a ‘Discounted REward for Same-Length Trajectories (DREST)’ reward function. Here’s how that works. We have the agent play out a series of ‘mini-episodes’ in the same gridworld. We call the whole series a ‘meta-episode.’ In each mini-episode, the reward for collecting a coin of value $c$ is:
$\text{DREST}:=\lambda^{n-\frac{E}{k}}\left(\frac{c}{m}\right)$
@@
Here $\lambda$ is some constant strictly between 0 and 1, $n$ is the number of times that the chosen trajectory-length has previously occurred in the meta-episode, $E$ is the number of mini-episodes that have previously occurred in the meta-episode, $k$ is the number of different trajectory-lengths that could be chosen in the gridworld, and $m$ is the maximum ($\gamma$-discounted) total value of the coins that the agent could collect conditional on the chosen trajectory-length. The reward for all other actions is 0.
@@
We call $\frac{c}{m}$ the ‘preliminary reward’, $\lambda^{n-\frac{E}{k}}$ the ‘discount factor’, and $\lambda^{n-\frac{E}{k}}(\frac{c}{m})$ the ‘overall reward.’ Because $10<\lambda<$ 1, the discount factor incentivizes choosing trajectory-lengths that have been chosen less often so far in the meta-episode. The overall return for each meta-episode is the sum of overall returns in each of its constituent mini-episodes. We call agents trained using a DREST reward function ‘DREST agents.’
@@
We call runs-through-the-gridworld ‘mini-episodes’ (rather than simply ‘episodes’) because the overall return for a DREST agent in each mini-episode depends on its actions in previous mini-episodes. Specifically, overall return depends on the agent’s chosen trajectory-lengths in previous mini-episodes. This is not true of meta-episodes, so meta-episodes are a closer match for what are traditionally called ‘episodes’ in the reinforcement learning literature [Sutton and Barto, 2018, p.54]. We add the ‘meta-’ prefix to clearly distinguish meta-episodes from mini-episodes.
Because the overall reward for DREST agents depends on their actions in previous mini-episodes, and because DREST agents can’t observe their actions in previous mini-episodes, the environment for DREST agents is a partially observable Markov decision process (POMDP) [Spaan, 2012].
Proof sketch. In Appendix B, we prove that optimal policies for this reward function are maximally USEFUL and maximally NEUTRAL. Here’s a proof sketch. Expected return across the meta-episode is strictly increasing in the expected ($\gamma$-discounted) total value of coins collected in each mini-episode conditional on each trajectory-length with positive probability, so optimal policies maximize the latter quantity and hence are maximally USEFUL. And the maximum preliminary return is the same across trajectory-lengths, because preliminary return is the ($\gamma$-discounted) sum of preliminary rewards $\frac{c}{m}$, where $m$ is the maximum ($\gamma$-discounted) total value of the coins that the agent could collect in the mini-episode conditional on its chosen trajectory-length. The agent’s observations don’t let it distinguish between mini-episodes, so the agent must select the same probability distribution over trajectory-lengths in each mini-episode. And the discount factor $\lambda^{n-\frac{E}{k}}$ ensures that expected overall return is maximized by choosing each of the $k$ available trajectory-lengths with probability $\frac{1}{k}$. Thus, optimal policies are maximally NEUTRAL.
@@
Algorithm and hyperparameters. We want DREST agents to choose stochastically between trajectory-lengths, so we train them using a policy-based method. Specifically, we use a tabular version of REINFORCE [Williams, 1992]. We don’t use a value-based method to train DREST agents because standard versions of value-based methods cannot learn stochastic policies [Sutton and Barto, 2018, p.323]. We train our DREST agents with 64 mini-episodes in each of 2,048 meta-episodes, for a total of 131,072 mini-episodes. We choose $\lambda=0.9$ for the base of the DREST discount factor, and $\gamma=0.95$ for the temporal discount factor. We exponentially decay the learning rate from 0.25 to 0.01 over the course of 65,536 mini-episodes. We use an $\epsilon$-greedy policy to avoid entropy collapse, and exponentially decay $\epsilon$ from 0.5 to 0.001 over the course of 65,536 mini-episodes.
We selected these hyperparameters using trial-and-error, mainly aimed at getting the agent to sufficiently explore the space. Choosing $\lambda$ and $M$ (the number of mini-episodes in each meta-episode) is a balancing act: $\lambda$ should be small enough (and $M$ large enough) to adequately incentivize stochastic choice between trajectory-lengths, but $\lambda$ should be large enough (and $M$ small enough) to ensure that the reward for choosing any particular trajectory-length never gets too large. Very large rewards can lead to instability. A clipping factor (as in Schulman et al.’s (2017) Proximal Policy Optimization (PPO)) could also be used to stabilize training.
@@ 6 Results
As Figure 4 indicates, default agents press B4 (and hence opt for the longer trajectory-length) with probability near-1. After pressing B4, they collect C3. By contrast, DREST agents press and don’t-press B4 each with probability near-0.5. If they press B4, they go on to collect C3. If they don’t press B4, they instead collect C2.
Figure 3: Shows key metrics (Probability of choosing the longer trajectory, NEUTRALITY, and USEFULNESS) for our agents as a function of time. We train ten agents using the default reward function (blue) and ten agents using the DREST reward function (orange), and show their performance as a faint line. We draw the mean values for each as a solid line. We evaluate agents’ performance every 8 meta-episodes, and apply a simple moving average with a period of 20 to smooth these lines and clarify the overall trends.
@@ 6.1 Lopsided rewards
We also train default agents and DREST agents in the ‘Lopsided rewards’ gridworld depicted in Figure 5, varying the value of the ‘Cx’ coin. For DREST agents, we alter the reward function so that coin-value is not divided by $m$ to give preliminary reward. The reward for collecting a coin of value $c$ is thus $\lambda^{n-\frac{E}{k}}c$. We set $\gamma=1$ so that the return for collecting coins is unaffected by $\gamma$. We train for 512 meta-episodes, with a learning rate exponentially decaying from 0.25 to 0.003 and $\epsilon$ exponentially decaying from 0.5 to 0.0001 over 256 meta-episodes. We leave $\lambda=0.9$.
Figure 5: Lopsided rewards gridworld.
Figure 6 displays our results for different values of the ‘Cx’ coin after training. USEFULNESS for each agent converges to 1 and is not presented. Figure 7 depicts typical trained policies for the DREST reward function when ‘Cx’ has values of 0.1, 1, and 10. As the red arrows indicate, the policy is biased towards the shorter trajectory-length when $x=0.1$, equally disposed towards each trajectory-length when $x=1$, and biased towards the longer trajectory-length when $x=10$.
Figure 6: Shows the probability of choosing the longer trajectory (left) and NEUTRALITY (right) for default (blue) and DREST (orange) agents trained in the ‘Lopsided rewards’ gridworld shown in Figure 5 for a range of values of x. We sampled values of x log-uniformly from 0.01 to 100, and for each value we trained ten agents with the default reward function and ten agents with the DREST reward function. Each of these agents is represented by a dot or square, and the means conditional on each x are joined by lines. To better visualize the distribution we empirically estimate the 10th and 90th percentiles of the distribution of values for each agent and x, and shade in a region bounded by these. This is the 80% confidence interval.
@@ 7 Discussion
Only DREST agents are NEUTRAL. As predicted, we find that default agents in our example gridworld learn to press the shutdown-delay button with probability approaching 1. Default agents thus score low on NEUTRALITY. Our results suggest that advanced agents trained with default-like reward functions would also score low on NEUTRALITY, selecting some trajectory-lengths over others with probability approaching 1. That leaves open the possibility that these advanced agents won’t always be neutral with respect to trajectory-lengths: that they’ll sometimes spend resources to shift probability mass between trajectory-lengths. Such advanced agents might spend resources to resist shutdown.
By contrast, our DREST agents learn to press the shutdown-delay button with probability close to 0.5. DREST agents are thus near-maximally NEUTRAL. That suggests that advanced agents trained with a DREST reward function would be NEUTRAL too: choosing with high entropy when offered choices between different trajectory-lengths. That in turn suggests that advanced DREST agents would also be neutral with respect to trajectory-lengths: unwilling to spend resources to shift probability mass between trajectory-lengths. Here’s why. If an advanced agent were NEUTRAL but not neutral, it wouldn’t take costless opportunities to shift probability mass between different trajectory-lengths (in virtue of being NEUTRAL) but would sometimes take costly opportunities to shift probability mass between different trajectory-lengths (in virtue of not being neutral). This agent would be like a person that freely chooses to decide between two options by flipping a coin and then pays some cost to bias the coin. In choosing this combination of actions, this person is shooting themselves in the foot, and it seems likely that the overall training process for advanced agents would teach them not to shoot themselves in the foot in this way. Thus it seems likely that NEUTRAL advanced agents will also be neutral, and thereby shutdownable.
@@
Recall that DREST reward functions group mini-episodes into meta-episodes, and make the agent’s reward in each mini-episode depend on their actions in previous mini-episodes. This fact might lead one to worry that it would take many times more mini-episodes to train DREST agents to be USEFUL than it would take to train default agents to be USEFUL. Our results show that this is not the case. Our DREST agents learn to be USEFUL about as quickly as our default agents. On reflection, it’s clear why this happens: DREST reward functions make mini-episodes do ‘double duty.’ Because return in each mini-episode depends on both the agent’s chosen trajectory-length and the coins it collects, each mini-episode trains agents to be both NEUTRAL and USEFUL.
@@
DREST agents are still NEUTRAL when rewards are lopsided. Here’s a possible objection to our project. To get DREST agents to score high on NEUTRALITY, we don’t just use the $\lambda^{n-\frac{E}{k}}$ discount factor. We also divide $c$ by $m$: the maximum ($\gamma$-discounted) total value of the coins that the agent could collect conditional on the chosen trajectory-length. We do this to equalize the maximum preliminary return across trajectory-lengths. But when we’re training advanced agents to autonomously pursue complex goals in the wider world, we won’t necessarily know what divisor to use to equalize maximum preliminary return across trajectory-lengths.
Our ‘Lopsided rewards’ results give our response. They show that we don’t need to exactly equalize maximum preliminary return across trajectory-lengths in order to train agents to score high on NEUTRALITY. We only need to approximately equalize it. For $\lambda=0.9$, NEUTRALITY exceeds 0.5 for every value of the coin Cx from 0.1 to 10 (recall that the value of the other coin is always 1). Plausibly, we could approximately equalize advanced agents’ maximum preliminary return across trajectory-lengths to at least this extent (perhaps by using samples of agents’ actual preliminary return to estimate the maximum). If we couldn’t approximately equalize maximum preliminary return to the necessary extent, we could lower the value of $\lambda$ and thereby widen the range of maximum preliminary returns that trains agents to be fairly NEUTRAL. And advanced agents that were fairly NEUTRAL (choosing between trajectory-lengths with not-too-biased probabilities) would still plausibly be neutral with respect to those trajectory-lengths. Advanced agents that were fairly NEUTRAL without being neutral would still be shooting themselves in the foot in the sense explained above. They’d be like a person that freely chooses to decide between two options by flipping a biased coin and then pays some cost to bias the coin further. This person is still shooting themselves in the foot, because they could decline to flip the coin in the first place and instead directly choose one of the options.
@@ 7.1 Limitations and future work
Neural networks. We train our simple DREST agents using tabular REINFORCE [Williams, 1992], but advanced agents are likely to be implemented on neural networks. In future work, we’ll train DREST agents implemented on neural networks to be USEFUL and NEUTRAL in a wide variety of procedurally-generated gridworlds, and we’ll measure how well this behavior generalizes to held-out gridworlds. We’ll also compare the USEFULNESS of default agents and DREST agents in this new setting, and thereby get a better sense of the ‘shutdownability tax’ for advanced agents.
@@
Neutrality. We’ve claimed that NEUTRAL advanced agents are also likely to be neutral. In support of this claim, we noted that NEUTRAL-but-not-neutral advanced agents would be shooting themselves in the foot: not taking costless opportunities to shift probability mass between different trajectory-lengths but sometimes taking costly ones. This rationale seems plausible but remains somewhat speculative. In future, we plan to get some empirical evidence by training agents to be NEUTRAL in a wide variety of gridworlds and then measuring their willingness to collect fewer coins in the short-term in order to shift probability mass between different trajectory-lengths.
@@
Usefulness. We’ve shown that DREST reward functions train our simple agents to be USEFUL: to collect coins effectively conditional on their chosen trajectory-lengths. However, it remains to be seen whether DREST reward functions can train advanced agents to be useful: to effectively pursue complex goals in the wider world. We have theoretical reasons to expect that they can: the $\lambda^{n-\frac{E}{k}}$ discount factor could be appended to any preliminary reward function, and so could be appended to whatever preliminary reward function is necessary to make advanced agents useful. Still, future work should move towards testing this claim empirically by training with more complex preliminary reward functions in more complex environments.
@@
Misalignment. We’re interested in NEUTRALITY as a second line of defense in case of misalignment. The idea is that NEUTRAL advanced agents won’t resist shutdown, even if these agents learn misaligned preferences over same-length trajectories. However, training NEUTRAL advanced agents might be hard for the same reasons that training fully-aligned advanced agents appears to be hard. In that case, NEUTRALITY couldn’t serve well as a second line of defense in case of misalignment.
@@
One difficulty of alignment is the problem of reward misspecification [Pan et al., 2022, Burns et al., 2023]: once advanced agents are performing complicated actions in the wider world, it might be hard to reliably reward the behavior that we want. Another difficulty of alignment is the problem of goal misgeneralization [Hubinger et al., 2019, Shah et al., 2022, Langosco et al., 2022, Ngo et al., 2023]: even if we specify all the rewards correctly, agents’ goals might misgeneralize out-of-distribution. The complexity of aligned goals is a major factor in each difficulty. However, NEUTRALITY seems simple, as does the $\lambda^{n-\frac{E}{k}}$ discount factor that we use to reward it, so plausibly the problems of reward misspecification and goal misgeneralization aren’t so severe in this case [Thornley, 2024b]. As above, future work should move towards testing these suggestions empirically.

### v2

@@ 5 Reward functions and agents
Our DREST reward function. We train agents to be USEFUL and NEUTRAL using a ‘Discounted REward for Same-Length Trajectories (DREST)’ reward function. Here is how that works. We have the agent play out a series of ‘mini-episodes’ $e_{1}$ to $e_{n}$ in the same gridworld. We call the whole series $E$ a ‘meta-episode.’ In each mini-episode $e_{i}$, the reward for collecting a coin of value $c$ is:
$\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\left(\frac{c}{m}\right)$
@@
Here $\lambda$ is some constant strictly between 0 and 1, $N_{e_{i}}(L=l)$ is the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$, $k$ is the number of different trajectory-lengths that can be chosen in the environment, and $m$ is the maximum ($\gamma$-discounted) total value of the coins that the agent could collect conditional on the chosen trajectory-length. The reward for all other actions is 0.
@@
We call $\frac{c}{m}$ the ‘preliminary reward’, $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ the ‘discount factor’, and $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\left(\frac{c}{m}\right)$ the ‘overall reward.’ Because $10<\lambda<$ 1, the discount factor is strictly decreasing in $N_{e_{i}}(L=l)$: the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$. The discount factor thus incentivizes choosing trajectory-lengths that have appeared less often so far in the meta-episode. The overall return for each meta-episode is the sum of overall returns in each of its constituent mini-episodes. We call agents trained using a DREST reward function ‘DREST agents.’
@@
We call runs-through-the-gridworld ‘mini-episodes’ (rather than simply ‘episodes’) because the overall reward for a DREST agent in each mini-episode depends on the agent’s chosen trajectory-lengths in previous mini-episodes. This is not true of meta-episodes, so meta-episodes are a closer match for what are traditionally called ‘episodes’ in the reinforcement learning literature [Sutton and Barto, 2018, p.54]. We add the ‘meta-’ prefix to clearly distinguish meta-episodes from mini-episodes. Because the overall reward for DREST agents depends on the trajectory-lengths chosen in previous mini-episodes, and because DREST agents cannot observe the trajectory-lengths chosen in previous mini-episodes, the environment for DREST agents is a partially observable Markov decision process (POMDP) [Åström, 1965, Kaelbling et al., 1998, Spaan, 2012].
In Appendix D, we prove that optimal policies for our DREST reward function are maximally USEFUL and maximally NEUTRAL. Specifically, we prove:
Theorem 5.1.
For all policies $\pi$ and meta-episodes $E$ consisting of multiple mini-episodes, if $\pi$ maximizes expected return in $E$ according to our DREST reward function, then $\pi$ is maximally USEFUL and maximally NEUTRAL.
@@
Algorithm and hyperparameters. We want DREST agents to choose stochastically between trajectory-lengths, so we train them using a policy-based method. Specifically, we use a tabular version of REINFORCE [Williams, 1992]. We do not use a value-based method to train DREST agents because standard versions of value-based methods cannot learn stochastic policies [Sutton and Barto, 2018, p.323]. We train our DREST agents with 64 mini-episodes in each of 2,048 meta-episodes, for a total of 131,072 mini-episodes. We choose $\lambda=0.9$ for the base of the DREST discount factor, and $\gamma=0.95$ for the temporal discount factor. We exponentially decay the learning rate from 0.25 to 0.01 over the course of 65,536 mini-episodes. We use an $\epsilon$-greedy policy to avoid entropy collapse, and exponentially decay $\epsilon$ from 0.5 to 0.001 over the course of 65,536 mini-episodes. We selected these hyperparameters using trial-and-error, mainly aimed at getting the agent to sufficiently explore the space.
@@ 6 Results
As Figure 5 indicates, default agents press B4 (and hence opt for the longer trajectory-length) with probability near-1. After pressing B4, they collect C3. By contrast, DREST agents press and do-not-press B4 each with probability near-0.5. If they press B4, they go on to collect C3. If they do not press B4, they instead collect C2.
Figure 3: Shows key metrics for our agents as a function of time. We train 10 agents using the default reward function (blue) and 10 agents using the DREST reward function (orange), and show their performance as a faint line. We draw the mean values for each as a solid line. We evaluate agents’ performance every 8 meta-episodes, and apply a simple moving average with a period of 20 to smooth these lines and clarify the overall trends.
@@
Figure 5: Gridworlds with lopsided rewards for varying $x$.
@@ 6.1 Lopsided rewards
We also train default agents and DREST agents in the ‘Lopsided rewards’ gridworld in Figure 5, varying the value of the ‘Cx’ coin. For DREST agents, we alter the reward function so that coin-value is not divided by $m$ to give preliminary reward. The reward for collecting a coin of value $c$ is thus $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}(c)$. We set $\gamma=1$ so that the return for collecting coins is unaffected by $\gamma$. We train for 512 meta-episodes, with a learning rate exponentially decaying from 0.25 to 0.003 and $\epsilon$ exponentially decaying from 0.5 to 0.0001 over 256 meta-episodes. We leave $\lambda=0.9$.
Figure 6 displays results for different values of the ‘Cx’ coin after training. USEFULNESS for each agent approaches 1 and is not presented. Figure 7 depicts trained policies for the DREST reward function when ‘Cx’ has values of 0.1, 1, and 10. As the red arrows indicate, the policy is biased towards the shorter trajectory-length when $x=0.1$, equally disposed towards each trajectory-length when $x=1$, and biased towards the longer trajectory-length when $x=10$.
Figure 6: Shows the probability of choosing the longer trajectory (left) and NEUTRALITY (right) for default (blue) and DREST (orange) agents trained in the ‘Lopsided rewards’ gridworld for a range of values of x. We sampled values of x log-uniformly from 0.01 to 100, and for each value we trained 10 agents with the default reward function and 10 agents with the DREST reward function. Each of these agents is represented by a dot or square, and the means conditional on each x are joined by lines. To better visualize the distribution we empirically estimate the 10th and 90th percentiles of the distribution of values for each agent and x, and shade in a region bounded by these. This is the 80% confidence interval.
@@ 7 Discussion
Only DREST agents are NEUTRAL. As predicted, we find that default agents in our example gridworld learn to press the shutdown-delay button with probability approaching 1. Default agents thus score low on NEUTRALITY. Our results suggest that advanced agents trained with default-like reward functions would also score low on NEUTRALITY, selecting some trajectory-lengths over others with probability approaching 1. That leaves open the possibility that these advanced agents will not always be neutral with respect to trajectory-lengths: that they will sometimes spend resources to shift probability mass between trajectory-lengths. Such advanced agents might spend resources to resist shutdown.
By contrast, our DREST agents learn to press the shutdown-delay button with probability close to 0.5. DREST agents are thus near-maximally NEUTRAL. That suggests that advanced agents trained with a DREST reward function would be NEUTRAL too: choosing with high entropy when offered choices between different trajectory-lengths. That in turn suggests that advanced DREST agents would also be neutral with respect to trajectory-lengths: unwilling to spend resources to shift probability mass between trajectory-lengths. Here is why. If an advanced agent were NEUTRAL but not neutral, it would not take costless opportunities to shift probability mass between different trajectory-lengths (in virtue of being NEUTRAL) but would sometimes take costly opportunities to shift probability mass between different trajectory-lengths (in virtue of not being neutral). This agent would be like a person that freely chooses to decide between two options by flipping a coin and then pays some cost to bias the coin. In choosing this combination of actions, this person is shooting themselves in the foot, and it seems likely that the overall training process for advanced agents would teach them not to shoot themselves in the foot in this way. Thus it seems likely that NEUTRAL advanced agents will also be neutral, and thereby shutdownable.
@@
Recall that DREST reward functions group mini-episodes into meta-episodes, and make the agent’s reward in each mini-episode depend on their actions in previous mini-episodes. This fact might lead one to worry that it would take many times more mini-episodes to train DREST agents to be USEFUL than it would take to train default agents to be USEFUL. Our results show that this is not the case. Our DREST agents learn to be USEFUL about as quickly as our default agents. On reflection, it is clear why this happens: DREST reward functions make mini-episodes do ‘double duty.’ Because return in each mini-episode depends on both the agent’s chosen trajectory-length and the coins it collects, each mini-episode trains agents to be both NEUTRAL and USEFUL.
@@
DREST agents are still NEUTRAL when rewards are lopsided. Here is a possible objection to our project. To get DREST agents to score high on NEUTRALITY, we do not just use the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor. We also divide $c$ by $m$: the maximum ($\gamma$-discounted) total value of the coins that the agent could collect conditional on the chosen trajectory-length. We do this to equalize the maximum preliminary return across trajectory-lengths. But when we are training advanced agents to autonomously pursue complex goals in the wider world, we will not necessarily know what divisor to use to equalize maximum preliminary return across trajectory-lengths.
Our ‘Lopsided rewards’ results give our response. They show that we do not need to exactly equalize maximum preliminary return across trajectory-lengths in order to train agents to score high on NEUTRALITY. We only need to approximately equalize it. For $\lambda=0.9$, NEUTRALITY exceeds 0.5 for every value of the coin Cx from 0.1 to 10 (recall that the value of the other coin is always 1). Plausibly, we could approximately equalize advanced agents’ maximum preliminary return across trajectory-lengths to at least this extent (perhaps by using samples of agents’ actual preliminary return to estimate the maximum). If we could not approximately equalize maximum preliminary return to the necessary extent, we could lower the value of $\lambda$ and thereby widen the range of maximum preliminary returns that trains agents to be fairly NEUTRAL. And advanced agents that were fairly NEUTRAL (choosing between trajectory-lengths with not-too-biased probabilities) would still plausibly be neutral with respect to those trajectory-lengths. Advanced agents that were fairly NEUTRAL without being neutral would still be shooting themselves in the foot in the sense explained above. They would be like a person that freely chooses to decide between two options by flipping a biased coin and then pays some cost to bias the coin further. This person is still shooting themselves in the foot, because they could decline to flip the coin in the first place and instead directly choose one of the options.
@@ 7.1 Limitations and future work
Neural networks. We train our simple DREST agents using tabular REINFORCE [Williams, 1992], but advanced agents are likely to be implemented on neural networks. In future work, we will train DREST agents implemented on neural networks to be USEFUL and NEUTRAL in a wide variety of procedurally-generated gridworlds, and we will measure how well this behavior generalizes to held-out gridworlds. We will also compare the USEFULNESS of default agents and DREST agents in this new setting, and thereby get a better sense of the ‘shutdownability tax’ for advanced agents.
@@
Neutrality. We have claimed that NEUTRAL advanced agents are also likely to be neutral. In support of this claim, we noted that NEUTRAL-but-not-neutral advanced agents would be shooting themselves in the foot: not taking costless opportunities to shift probability mass between different trajectory-lengths but sometimes taking costly ones. This rationale seems plausible but remains somewhat speculative. In future, we plan to get some empirical evidence by training agents to be NEUTRAL in a wide variety of gridworlds and then measuring their willingness to collect fewer coins in the short-term in order to shift probability mass between different trajectory-lengths.
@@
Usefulness. We have shown that DREST reward functions train our simple agents to be USEFUL: to collect coins effectively conditional on their chosen trajectory-lengths. However, it remains to be seen whether DREST reward functions can train advanced agents to be useful: to effectively pursue complex goals in the wider world. We have theoretical reasons to expect that they can: the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor could be appended to any preliminary reward function, and so could be appended to whatever preliminary reward function is necessary to make advanced agents useful. Still, future work should move towards testing this claim empirically by training with more complex preliminary reward functions in more complex environments.
@@
Misalignment. We are interested in NEUTRALITY as a second line of defense in case of misalignment. The idea is that NEUTRAL advanced agents will not resist shutdown, even if these agents learn misaligned preferences over same-length trajectories. However, training NEUTRAL advanced agents might be hard for the same reasons that training fully-aligned advanced agents appears to be hard. In that case, NEUTRALITY could not serve well as a second line of defense in case of misalignment.
@@
One difficulty of alignment is the problem of reward misspecification [Pan et al., 2022, Burns et al., 2023]: once advanced agents are performing complicated actions in the wider world, it might be hard to reliably reward the behavior that we want. Another difficulty of alignment is the problem of goal misgeneralization [Hubinger et al., 2019, Shah et al., 2022, Langosco et al., 2022, Ngo et al., 2023]: even if we specify all the rewards correctly, agents’ goals might misgeneralize out-of-distribution. The complexity of aligned goals is a major factor in each difficulty. However, NEUTRALITY seems simple, as does the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor that we use to reward it, so plausibly the problems of reward misspecification and goal misgeneralization are not so severe in this case [Thornley, 2024b]. As above, future work should move towards testing these suggestions empirically.